please help me solve this equation:
solve the equation. check your answer.
5 - w/w + 5/2 = -4/w

Answer by jorel1380(3719)

5 - w/w + 5/2 = -4/w
5-1+5/2=-4/w
13/2=-4/w
13w=-8
w=-8/13..